OpenStudy (anonymous):

find an equation of variation in which y varies as x and the following is true y=40 and x=38

myininaya (myininaya):

y=kx

myininaya (myininaya):

directly right?

OpenStudy (dumbcow):

40 = 38k k = 40/38 = 20/19 y = (20/19)x

OpenStudy (anonymous):

no i no its 1520 just dont no what form to put it in

OpenStudy (anonymous):

my choices are y=1520 y=1520/x y=1.05/x y=1.05x

OpenStudy (amistre64):

if those are your choices then y = 1.05 x is your best shot

OpenStudy (dumbcow):

i gave you the answer in fraction form, given a fraction, put in your calculator as 20 divided by 19

OpenStudy (anonymous):

mr.cow your right i just divided how did u get the fraction i got 1.05

OpenStudy (dumbcow):

plug in x and y into general formula for direct variation y=kx solve for k 40 = 38k k = 40/38 = 20/19 y = (20/19)x

OpenStudy (anonymous):

thanxs u 2